# Recovery of Sparse Graph Signals

### v1

@@ Abstract
This paper investigates the recovery of a node-domain sparse graph signal from the output of a graph filter. This problem, which is often referred to as the identification of the source of a diffused sparse graph signal, is seminal in the field of graph signal processing (GSP). Sparse graph signals can be used in the modeling of a variety of real-world applications in networks, such as social, biological, and power systems, and enable various GSP tasks, such as graph signal reconstruction, blind deconvolution, and sampling. In this paper, we assume double sparsity of both the graph signal and the graph topology, as well as a low-order graph filter. We propose three algorithms to reconstruct the support set of the input sparse graph signal from the graph filter output samples, leveraging these assumptions and the generalized information criterion (GIC). First, we describe the graph multiple GIC (GM-GIC) method, which is based on partitioning the dictionary elements (graph filter matrix columns) that capture information on the signal into smaller subsets. Then, the local GICs are computed for each subset and aggregated to make a global decision. Second, inspired by the well-known branch and bound (BNB) approach, we develop the graph-based branch and bound GIC (graph-BNB-GIC), and incorporate a new tractable heuristic bound tailored to the graph and graph filter characteristics. Finally, we propose the graph-based first order correction (GFOC) method, which improves existing sparse recovery methods by iteratively examining potential improvements to the GIC cost function through replacing elements from the estimated support set with elements from their one-hop neighborhood. We conduct simulations over the square grid graph and the Erdős-Rényi graph that demonstrate that the proposed sparse recovery methods outperform the orthogonal matching pursuit (OMP) method, the Lasso method, and the BNB method with $\ell_{1}$ regularization in terms of support set recovery accuracy, and without a significant computational overhead.
@@ I Introduction
gsp theory extends classical time-varying signals to irregular domains represented by graphs [2, 3, 4]. The graph under consideration may represent a physical network, such as electrical or sensor networks, or a virtual network, such as social networks, or it may encode statistical dependencies among the signal values. Recent developments in graph signal processing (GSP) encompass a wide array of processing tools, including graph spectral analysis [5, 6], anomaly detection [7], sampling and signal recovery [8, 9, 10, 11, 12, 13, 14], verification of graph smoothness [15, 16], and graph filter design [2, 3, 17, 18]. However, limited attention has been paid to developing GSP-based techniques for the sparse recovery of node-domain sparse graph signals.
@@
Signal modeling based on sparse representations is used in numerous signal and image processing applications [19, 20, 21, 22]. Sparse recovery methods can be categorized into the following main groups: convex relaxations techniques, often using methods such as $\ell_{1}$-norm minimization or basis pursuit [19, 23, 24, 25, 26], non-convex optimization techniques [27, 28, 29], and greedy algorithms [30, 20, 31, 32, 33]. Bayesian inference approaches estimate the sparse signal by incorporating prior information, whereas other methods use knowledge of dictionary matrix structure. In network science, compressed sensing and group testing have been discussed in [34, 35, 36], while diffusion processess have been discussed in [37, 38, 39, 40, 41]. Additionally, several GSP works have investigated the reconstruction of bandlimited graph signals where the measurement vector belongs to a subspace defined by a small number of graph frequencies [42, 43, 44].
@@
Recently there has been growing interest in setups where a node-domain sparse graph signal ${{\bf{x}}}$ is filtered by a graph filter ${\bf{H}}$. In this case, the output signal ${{\bf{y}}}$ is referred to as a diffused sparse graph signal [45]. Modeling diffusion processes over graphs by graph filters is valuable due to the ability to capture local interactions among nodes [46, 47, 48, 49]. Diffused sparse graph signals can be used for modeling many real-world network scenarios. For instance, in social networks, these signals can be used for opinion forming [50] and finding the origin of a rumor [51]. In biological networks, they enable the analysis of the spread of a disease [52], and in computer networks, they can be used to locate malware [53]. In power systems, they may be used for locating anomalies in power signals [54, 55]. In the field of GSP, diffused sparse graph signals have been studied in several problems, such as bandlimited graph signal reconstruction [56]. Blind deconvolution, in which both the graph filter and the sparse input are estimated, has been discussed in [45, 46, 57, 58]. System identification, considered in [47, 45], refers to the case where the sparse signal is known, but the graph filter coefficients are estimated. The sampling of diffused sparse graph signals is discussed in [59]. However, in the applications discussed above, the sparse recovery does not exploit the GSP-based modeling of the sparse graph signals. Instead, the sparse recovery is typically done by sparse approximations, usually $\ell_{1}$ relaxation, or with a tailored solution that uses the specific properties of the application.
@@
This paper establishes sparse recovery methods for graph signals using the generalized information criterion (GIC) as the cost function, and practical assumptions including the sparsity of the underlying graph. A direct solution for the GIC-based problem necessitates an exhaustive search, whose complexity grows exponentially with the size of the graph, rendering it impractical. We propose three graph-based sparse recovery methods that harvest the double sparsity imposed by the graph and the graph signal. The first method, the graph-based multiple generalized information criterion (GM-GIC) method, identifies and partitions a set of suspected nodes and then applies a local GIC on each partition, which significantly reduces the computational overhead of the GIC-based approach. The second method, the graph-based branch and bound GIC (Graph-BNB-GIC) method, further mitigates the computational demands by iteratively searching over the candidate support sets of the sparse graph signals (the same sets as in the GM-GIC method). This method is developed based on the branch and bound (BNB) method [60, 61] with a heuristic graph-based upper bound that exploits the double sparsity and the graph properties. The third method, the graph-based first order correction (GFOC) method, builds upon existing sparse recovery techniques and improves them in terms of the GIC cost function by replacing elements from the estimated support set with elements from their one-hop neighborhood. Our simulations over the square grid graph and an Erdős-Rényi graph demonstrate that the proposed methods outperform the state-of-the-art orthogonal matching pursuit (OMP), Lasso, and BNB with $\ell_{1}$ regularization methods in terms of the support set recovery accuracy, and without a significant computation overhead.
@@
The paper is organized as follows. Section II presents the necessary background and outlines the problem formulation. In Section III, we analyze the dictionary matrix for the proposed graphical setting. We describe both the GM-GIC and Graph-BNB-GIC methods in Section IV, followed by the presentation of the GFOC method in Section V. In Section VI, we present the simulation results, and the conclusions are provided in Section VII.
@@ II-A Background: Graphs and Graph Filters
A graph shift operator (GSO), ${{{\bf{S}}}\in\mathbb{R}^{{\lvert{\mathcal{V}}\rvert}\times{\lvert{\mathcal{V}}\rvert}}}$, is a linear operator applied on graph signals that satisfies
@@
where the geodesic distance between nodes $k$ and $m$, $\delta(k,m)$, is the number of edges on the shortest path between these nodes. According to (3), the non-zero elements of ${{{{\bf{S}}}}_{k,m}}$ are the diagonal elements (i.e. $k=m$) or the elements that are associated with an edge (i.e. ${(k,m)}\in{\mathcal{E}}$). As a result, by applying a graph shift operator on a graph signal ${{\bf{x}}}$, each element of the shifted signal, ${{\bf{S}}}{{\bf{x}}}$, is a linear combination of the signal elements in its one-hop neighborhood:
@@
Thus, the transformation by the graph shift operator can be computed locally at each node by aggregating the values of the input signal within the one-hop neighborhood of each node.
A shift-invariant graph filter, ${{{\bf{H}}}\in\mathbb{R}^{{\lvert{\mathcal{V}}\rvert}\times{\lvert{\mathcal{V}}\rvert}}}$, is defined as a polynomial of the graph shift operator [3, 4]:
@@
where $h_{0},\ldots,h_{{\Psi}}$ are the coefficients of the graph filter and $1\leq{\Psi}\leq{\lvert{\mathcal{V}}\rvert}$ is the graph filter order. By substituting (3) in (6) it can be verified that the $(k,m)$th element of ${{\bf{H}}}$ satisfies
@@ II-B Assumptions on the Graphical Setting
    The graph filter is a polynomial of the graph shift operator as defined in (6). Aditionally, the graph filter degree, ${\Psi}$, is small with respect to (w.r.t.) the number of nodes, i.e. ${\Psi}\ll{\lvert{\mathcal{V}}\rvert}$.
@@
The graph filter in (6) under Assumptions A.1-A.2 can be used to model a variety of GSP tasks, including signal reconstruction, blind deconvolution, sampling and system identification [45, 46, 57, 47, 59]. Additionally, this framework can be used for modeling real-life applications in social networks [50, 51], in biological networks [52], in computer networks [53], in wireless sensor networks [5], and in power systems [54, 55]. It is worth noting that in the context of power systems, the average nodal degree is generally very low (usually between 2 to 5) and does not increase with the network size [62]. This property is also common in other infrastructure systems, such as water and transportation systems.
@@ II-C Recovery of Sparse Graph Signals
The inner minimization problem in (14) is a least squares (LS) problem with the solution (see, e.g. p. 225 in [63]):
@@
The support set recovery problem presented in (16) qualifies for any sparse signal ${{\bf{x}}}$ and dictionary matrix ${{\bf{H}}}$, not necessarily over graphs, and is valid as long as ${{\bf{H}}}_{{\Omega}}^{T}{{\bf{H}}}_{{\Omega}}^{\leavevmode\nobreak\ }$ is a nonsingular matrix. This problem can be viewed as a multiple-hypothesis testing problem, where under each hypothesis, the sparse signal has a different support set [65], which belongs to the following set of candidate sets:
@@
A direct solution of the support set recovery in (16) tends to overfit by selecting support sets with larger cardinality than the true support set. To mitigate this problem, we use the well-known GIC model selection method, which is a generalization of various criteria, such as the Akaike information criterion (AIC) and the minimum description length (MDL). For the considered problem, the GIC is given by [66]
$\hat{{\Omega}}=\arg\underset{{\lvert{\Omega}\rvert}\leq{s}}{\max}{\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}}}}}{{\bf{y}}}\|}^{2}-{\rho({\lvert{\Omega}\rvert})}.$ (18)
Compared to (16), the GIC in (18) includes an additional penalty function component, ${\rho({\lvert{\Omega}\rvert})}$, which depends on the support set cardinality and mitigates the risk of overfitting. After solving (18), the sparse signal recovery solution can be obtained by substituting $\hat{{\Omega}}$ in (15).
@@ III Dictionary Matrix Analysis
where ${{\bf{H}}}_{k}$ is the $k$th column of ${{\bf{H}}}$. Hence, the filtered signal, ${{\bf{H}}}{{\bf{x}}}$, can be viewed as a linear combination of dictionary atoms (graph filter columns). In addition, if two dictionary atoms correspond to nodes $k$ and $m$ with a geodesic distance larger than $2{\Psi}$, i.e. $\delta(k,m)>{\Psi}$, then the projection matrix definition and (9) imply that
@@
where ${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}$ is the projection matrix onto the single-column space $\text{col}({{\bf{H}}}_{m})$ (see the definition in (1)).
Now, let ${\mathcal{N}_{\scriptstyle{2{\Psi}}}[{\Omega}_{t}]}$ be the ${\Psi}$-order neighborhood of the set ${\Omega}_{t}$. Hence,
${\mathcal{N}_{\scriptstyle{2{\Psi}}}[{\Omega}_{t}]}\stackrel{{\scriptstyle\triangle}}{{=}}\bigcup_{k\in{\Omega}_{t}}{\mathcal{N}_{\scriptstyle{2{\Psi}}}[k]}\hskip 99.58464pt$
@@
where ${\mathcal{N}_{\scriptstyle{2{\Psi}}}[k]}$ is defined in (5). Thus, assume that $m\notin{\mathcal{N}_{\scriptstyle{2{\Psi}}}[k]}$, then, left-multiplying (19) by ${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}$ and substituting (20) results in
${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{H}}}{{\bf{x}}}={{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{H}}}_{{\Omega}_{t}}{{\bf{x}}}_{{\Omega}_{t}}={\bf{0}},\leavevmode\nobreak\ \forall m\notin{\mathcal{N}_{\scriptstyle{2{\Psi}}}[{\Omega}_{t}]}.$ (22)
Therefore, left-multiplying the measurement model in (11) by ${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}$ and substituting (22) results in
${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{y}}}={{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{H}}}{{\bf{x}}}+{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{n}}}={{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{n}}},$ (23)
for all $m\notin{\mathcal{N}_{\scriptstyle{2{\Psi}}}[{\Omega}_{t}]}$. Consequently, the dictionary atoms, indexed by nodes outside the set ${\mathcal{N}_{\scriptstyle{2{\Psi}}}[{\Omega}_{t}]}$ in (III), contain only noise. That is, the dictionary matrix elements that contain information on the sparse graph signal are only those included in the $2{\Psi}$ neighborhood of ${\Omega}_{t}$, $\mathcal{N}_{2{\Psi}}[{\Omega}_{t}]$.
The following theorem establishes a bound on the cardinality of the set ${\mathcal{N}_{\scriptstyle{2{\Psi}}}[{\Omega}_{t}]}$ from (III) as a function of the graph filter degree ${\Psi}$, the graph structure, and the sparsity level $s$. This theorem is important as it provides constraints on the size of a reduced set of dictionary atoms containing signal information. It offers two key advantages: first, it reduces the number of candidate support sets used to construct the feasible set in (18); second, observing a reduced set of dictionary atoms enhances processing flexibility, where a smaller set may be viewed as multiple (unrelated) regions across the graph, while the entire node set, ${\mathcal{V}}$, must be regarded as a single region.
@@
Therefore, since ${\lvert{\Omega}_{t}\rvert}\leq{s}$ under the sparsity assumption in (12), we conclude that the $2{\Psi}$-neighborhood of ${\Omega}$ is bounded by the right hand side (r.h.s.) of (24). ∎
The bound in (24) depends on the sparsity level, ${s}$, the maximal degree of the graph, $d_{\max}$, and the graph filter degree, ${\Psi}$, that are all considered to be small based on (12), Assumption A.1, and Assumption A.2, respectively. In contrast, the bound is independent of the size of the graph, i.e. the number of nodes, ${\lvert{\mathcal{V}}\rvert}$. Thus, for large networks it can be assumed that the bound is significantly lower than the actual number of nodes:

### v2

# Efficient Recovery of Sparse Graph Signals from Graph Filter Outputs

@@ Abstract
This paper investigates the recovery of a node-domain sparse graph signal from the output of a graph filter. This problem, which is often referred to as the identification of the source of a diffused sparse graph signal, is seminal in the field of graph signal processing (GSP). Sparse graph signals can be used in the modeling of a variety of real-world applications in networks, such as social, biological, and power systems, and enable various GSP tasks, such as graph signal reconstruction, blind deconvolution, and sampling. In this paper, we assume double sparsity of both the graph signal and the graph topology, as well as a low-order graph filter. We propose three algorithms to reconstruct the support set of the input sparse graph signal from the graph filter output samples, leveraging these assumptions and the generalized information criterion (GIC). First, we describe the graph multiple GIC (GM-GIC) method, which is based on partitioning the dictionary elements (graph filter matrix columns) that capture information on the signal into smaller subsets. Then, the local GICs are computed for each subset and aggregated to make a global decision. Second, inspired by the well-known branch and bound (BNB) approach, we develop the graph-based branch and bound GIC (graph-BNB-GIC), and incorporate a new tractable heuristic bound tailored to the graph and graph filter characteristics. In addition, we propose the graph-based first order correction (GFOC) method, which improves existing sparse recovery methods by iteratively examining potential improvements to the GIC cost function by replacing elements from the estimated support set with elements from their one-hop neighborhood. Simulations on stochastic block model (SBM) graphs demonstrate that the proposed sparse recovery methods outperform existing techniques in terms of support set recovery and mean-squared-error (MSE), without significant computational overhead. In addition, we investigate the application of our graph-based sparse recovery methods in blind deconvolution scenarios where the graph filter is unknown. Simulations using real-world data from brain networks and pandemic diffusion analysis further demonstrate the superiority of our approach compared to graph blind deconvolution techniques.
@@ I Introduction
gsp theory extends classical time-varying signals to irregular domains represented by graphs [2, 3, 4]. The graph under consideration may represent a physical network, such as electrical or sensor networks, or a virtual network, such as social networks, or it may encode statistical dependencies among the signal values. Recent developments in gsp (gsp) encompass a wide array of processing tools, including graph spectral analysis [5, 6], anomaly detection [7], sampling and signal recovery [8, 9, 10, 11, 12, 13, 14], verification of graph smoothness [15, 16], and graph filter design [2, 3, 17, 18]. However, limited attention has been paid to developing gsp-based techniques for the sparse recovery of node-domain sparse graph signals.
@@
Signal modeling based on sparse representations is used in numerous signal and image processing applications [19, 20, 21, 22]. Sparse recovery methods can broadly be categorized into three main groups: convex relaxations, often using methods such as $\ell_{1}$-norm minimization or basis pursuit [19, 23, 24, 25, 26], non-convex optimization approaches [27, 28, 29], and greedy algorithms [30, 20, 31, 32, 33]. In the context of estimation theory, Bayesian inference approaches estimate the sparse signal by incorporating prior information, while other methods use the structure of the dictionary matrix. In network science, compressed sensing and group testing have been discussed in [34, 35, 36], and diffusion processes that originate from a sparse signal have been discussed in [37, 38, 39, 40, 41]. Several gsp studies have investigated the sampling and reconstruction of graph signals by assuming the signal belongs to a vector subspace. For example, the works [42, 43, 44, 45] assume graph bandlimited signals, while [46] considers sparsity in the spectral domain. Similarly, [47] models the graph signals as smooth in the node domain or decaying in the spectral domain.
@@
Recently there has been growing interest in setups where a node-domain sparse graph signal ${{\bf{x}}}$ is filtered by a graph filter ${\bf{H}}$. In this case, the output signal ${{\bf{y}}}$ is referred to as a diffused sparse graph signal [48]. Modeling diffusion processes over graphs by graph filters is valuable due to the ability to capture local interactions among nodes [49, 50, 51, 52]. Diffused sparse graph signals can be used for modeling many real-world network scenarios. For instance, in social networks, these signals can be used for opinion forming [53] and finding the origin of a rumor [54]. In biological networks, they enable the analysis of the spread of a disease [55], and in computer networks, they can be used to locate malware [56]. In power systems, they may be used for locating anomalies in power signals [57, 58]. In the field of gsp, diffused sparse graph signals have been studied in several problems, such as bandlimited graph signal reconstruction [59]. Blind deconvolution, in which both the graph filter and the sparse input are estimated, has been discussed in [48, 49, 60, 61, 62]. System identification, considered in [50, 48], refers to the case where the sparse signal is known, but the graph filter coefficients are estimated. The sampling of diffused sparse graph signals is discussed in [63]. For an overview of blind deconvolution and signal reconstruction methods for diffused sparse graph signals, we refer the readers to Fig. 1 in [48]. However, in most of the applications discussed above, the sparse recovery does not exploit the gsp-based modeling of the sparse graph signals. Instead, these methods typically rely on standard sparse approximations, such as $\ell_{1}$ relaxation, or employ tailored solutions that use specific properties of the application at hand.
@@
In this paper, we establish efficient sparse recovery methods for graph signals using the gic (gic) [64] as the cost function. We assume double sparsity of both the graph and the graph signal and that the network diffusion process is modeled by a known low-degree graph filter. A direct solution for the gic-based problem necessitates an exhaustive search, whose complexity grows exponentially with the size of the graph, rendering it impractical. We propose three graph-based sparse recovery methods that harvest the double sparsity imposed by the graph and the graph signal. The first method, the gmgic (gmgic) method, identifies and partitions a set of suspected nodes and then applies a local gic on each partition, which significantly reduces the computational overhead of the gic-based approach. The second method, the gbnb (gbnb) method, further mitigates the computational demands by iteratively searching over the candidate support sets of the sparse graph signals (the same sets as in the gmgic method). This method is developed based on the bnb (bnb) method [65, 66] with a heuristic graph-based upper bound that exploits the double sparsity and the graph properties. The third method, the gfoc (gfoc) method, builds upon existing sparse recovery techniques and improves them in terms of the GIC cost function by replacing elements from the estimated support set with elements from their one-hop neighborhood. Our simulations on sbm (sbm) graphs demonstrate that the proposed methods outperform state-of-the-art techniques: omp (omp), Lasso, and bnb with $\ell_{1}$ regularization methods in terms of the support set recovery accuracy and mse (mse), without a significant computation overhead. In addition, to address scenarios where the graph filter is unknown, we introduce a new approach that incorporates a preliminary blind deconvolution stage to estimate the graph filter. Our simulations demonstrate that the proposed graph-based recovery methods can significantly improve the performance of existing graph blind deconvolution techniques in real-world applications, including: (1) brain networks obtained from real data [67]; and (2) the identification of the source of a disease using data from the $1854$ cholera outbreak [68, 69].
@@
The paper is organized as follows. Section II presents the necessary background and outlines the problem formulation. In Section III, we analyze the dictionary matrix for the proposed graphical setting. We describe both the gmgic and gbnb methods in Section IV, followed by the presentation of the gfoc method in Section V. In Section VI, we present the simulation results, and the conclusions are provided in Section VII.
@@ II-A Background: Graphs and Graph Filters
A gso (gso), ${{{\bf{S}}}\in\mathbb{R}^{{\lvert{\mathcal{V}}\rvert}\times{\lvert{\mathcal{V}}\rvert}}}$, is a linear operator applied on graph signals that satisfies
@@
where the geodesic distance between nodes $k$ and $m$, $\delta(k,m)$, is the number of edges on the shortest path between these nodes. According to (3), the non-zero elements of ${{{{\bf{S}}}}_{k,m}}$ are the diagonal elements (i.e. $k=m$) or the elements that are associated with an edge (i.e. ${(k,m)}\in{\mathcal{E}}$). As a result, by applying a gso on a graph signal ${{\bf{x}}}$, each element of the shifted signal, ${{\bf{S}}}{{\bf{x}}}$, is a linear combination of the signal elements in its one-hop neighborhood:
@@
Thus, the transformation by the gso can be computed locally at each node by aggregating the values of the input signal within the one-hop neighborhood of each node.
A shift-invariant graph filter, ${{{\bf{H}}}\in\mathbb{R}^{{\lvert{\mathcal{V}}\rvert}\times{\lvert{\mathcal{V}}\rvert}}}$, is defined as a polynomial of the gso[3, 4]:
@@
where $h_{0},\ldots,h_{{\Psi}}$ are the coefficients of the graph filter and $1\leq{\Psi}\leq{\lvert{\mathcal{V}}\rvert}-1\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$ is the graph filter order. By substituting (3) in (6) it can be verified that the $(k,m)$th element of ${{\bf{H}}}$ satisfies
@@ II-B Assumptions on the Graphical Setting
    The graph filter is a polynomial of the gso as defined in (6). Additionally, the graph filter degree, ${\Psi}$, is small wrt (wrt) the number of nodes, i.e. ${\Psi}\ll{\lvert{\mathcal{V}}\rvert}$.
@@
The graph filter in (6) under Assumptions A.1-A.2, in addition to the sparsity restriction in (12), can be used to model a variety of gsp tasks, including signal reconstruction, blind deconvolution, sampling, and system identification [48, 49, 60, 50, 63]. Additionally, this framework can be used for modeling real-life applications in social networks [53, 54], biological networks [55], brain networks [67], epidemiology, computer networks [56], wireless sensor networks [5], and power systems [57, 58]. It is worth noting that in the context of power systems, the average nodal degree is generally very low (usually between 2 to 5) and does not increase with the network size [70]. This property is also common in other infrastructure systems, such as water and transportation systems.
@@ II-C Recovery of Sparse Graph Signals
The inner minimization problem in (14) is a ls (ls) problem with the solution (see, e.g. p. 225 in [71]):
@@
The support set recovery problem presented in (16) applies to any sparse signal ${{\bf{x}}}$ and dictionary matrix ${{\bf{H}}}$, and is not restricted to graph-based settings. This formulation is valid as long as ${{\bf{H}}}_{{\Omega}}^{T}{{\bf{H}}}_{{\Omega}}^{\leavevmode\nobreak\ }$ is a nonsingular matrix. Therefore, we assume that for any support set $\Omega$ that satisfies ${\lvert{\Omega}\rvert}\leq s$, i.e. support that is relevant to our algorithms, the matrix ${{\bf{H}}}_{{\Omega}}$ has a full column rank. This problem can be viewed as a multiple-hypothesis testing problem, where under each hypothesis, the sparse signal has a different support set [73], which belongs to the following set of candidate sets:
@@
A direct solution of the support set recovery in (16) tends to overfit by selecting support sets with larger cardinality than the true support set. To mitigate this problem, we use the well-known gic model selection method, which is a generalization of various criteria, such as the aic (aic) and the mdl (mdl). For the considered problem, the gic is given by [64]
$\hat{{\Omega}}=\arg\underset{{\lvert{\Omega}\rvert}\leq{s}}{\max}{\|{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{{\Omega}}}}}{{\bf{y}}}\|}^{2}-{\rho{\lvert{\Omega}\rvert}}.$ (18)
Compared to (16), the gic in (18) includes an additional penalty function component, ${\rho{\lvert{\Omega}\rvert}}$, which depends on the support set cardinality. This penalty term mitigates the risk of overfitting; a larger penalty will lead to less complex models with a smaller support set, while a smaller penalty will result in a larger support set [64]. The penalty function can be tuned based on historical data, through trial and error, or by using model prior information. After solving (18), the sparse signal recovery solution can be obtained by substituting $\hat{{\Omega}}$ in (15).
@@ III Dictionary Matrix Analysis
where ${{\bf{H}}}_{k}$ is the $k$th column of ${{\bf{H}}}$. Hence, the filtered signal, ${{\bf{H}}}{{\bf{x}}}$, can be viewed as a linear combination of dictionary atoms (graph filter columns). Therefore, projecting the filtered signal into the single-column space $\text{col}({{\bf{H}}}_{m})$ results in
${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{H}}}{\bf{x}}=\sum_{k\in{\Omega}_{t}}{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{H}}}_{k}{x}_{k},$ (20)
where ${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}$ (defined in (1)) is a linear operator. Thus, it can be verified that
${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{H}}}{\bf{x}}={\bf{0}}\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ {{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{H}}}_{k}={\bf{0}},\leavevmode\nobreak\ \forall k\in{\Omega}_{t}.$ (21)
In the following lemma, we show that ${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{H}}}_{k}=0$ if the geodesic distance between $k$ and $m$ is larger than $2{\Psi}$, $\delta(k,m)>2{\Psi}$.
Lemma 2.
If two dictionary atoms correspond to nodes $k$ and $m$ with a geodesic distance larger than $2{\Psi}$, i.e. $\delta(k,m)>2{\Psi}$, then
@@
where ${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}$ is the projection matrix onto the single-column space $\text{col}({{\bf{H}}}_{m})$.
Proof:
The first equality in (22) is obtained by substituting the definition of the projection matrix ${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}$. The second equality is obtained by substituting (9) from Lemma 1, since it is given that $m$ and $k$ satisfy $\delta(m,k)>2{\Psi}$. ∎
Leveraging the results in (19)-(22), we present the following definition and theorem.
Definition 1.
The ${\Psi}$-order neighborhood of the set ${\Omega}_{t}$ is
${\mathcal{N}_{\scriptstyle{2{\Psi}}}[{\Omega}_{t}]}\mathrel{{}\mathop{=}\limits^{\triangle}}\bigcup_{k\in{\Omega}_{t}}{\mathcal{N}_{\scriptstyle{2{\Psi}}}[k]}\hskip 123.76965pt$
@@
where ${\mathcal{N}_{\scriptstyle{2{\Psi}}}[k]}$ is defined in (5).
Theorem 3.
If $m\notin{\mathcal{N}_{\scriptstyle{2{\Psi}}}[k]}$, then the following hold:
${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{H}}}{{\bf{x}}}={\bf{0}},$ (24a)
${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{y}}}={{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{H}}}{{\bf{x}}}+{{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{n}}}={{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{n}}},$ (24b)
where ${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}$ is defined in (1) and ${{\bf{y}}}$ is the measurement model defined in (11).
Proof:
Since $m\notin{\mathcal{N}_{\scriptstyle{2{\Psi}}}[k]}$, then Lemma 2 implies that ${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{H}}}{{\bf{x}}}$ can be expressed using (20), where each of the components in the summation satisfies ${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}{{\bf{H}}}_{k}x_{k}={\bf{0}}$. Thus, we obtain (24a). Additionally, by left-multiplying the measurement model in (11) by ${{\bf{P}}_{\scriptstyle{{{\bf{H}}}_{m}}}}$ and then substituting (24a) in the result, we obtain (24b). ∎
It can be seen that Theorem 3 implies that the dictionary atoms indexed by nodes outside the set ${\mathcal{N}_{\scriptstyle{2{\Psi}}}[{\Omega}_{t}]}$ in (1) contain only the contribution of the noise. That is, the dictionary matrix elements that contain information on the sparse graph signal are only those included in the $2{\Psi}$ neighborhood of ${\Omega}_{t}$, $\mathcal{N}_{2{\Psi}}[{\Omega}_{t}]$. Thus, in practice, the true support of the signal is included in the set
$\{{\Omega}\in\mathcal{N}_{2{\Psi}}[{\Omega}_{t}]:\leavevmode\nobreak\ \lvert{\Omega}\rvert\leq s\}.$ (25)
Consequently, we can significantly reduce the search set in (18) by using (25) instead of (17). This point is formalized in the following theorem, which establishes a bound on the cardinality of the set ${\mathcal{N}_{\scriptstyle{2{\Psi}}}[{\Omega}_{t}]}$ from (1) as a function of the graph filter degree ${\Psi}$, the graph structure, and the sparsity level $s$. This theorem provides constraints on the size of a reduced set of dictionary atoms that contain signal information. In particular, it offers two key advantages: first, it reduces the number of candidate support sets used to construct the feasible set in (18); second, in some cases, the node set, ${\mathcal{V}}$, could be divided into smaller subsets, enabling parallel processing.
@@
Therefore, since ${\lvert{\Omega}_{t}\rvert}\leq{s}$ under the sparsity assumption in (12), we conclude that the $2{\Psi}$-neighborhood of ${\Omega}$ is bounded by the rhs (rhs) of (26). ∎
The bound in (26) depends on the sparsity level, ${s}$, the maximal degree of the graph, $d_{\max}$, and the graph filter degree, ${\Psi}$, which are all considered to be small based on (12), Assumption A.1, and Assumption A.2, respectively. In contrast, the bound is independent of the size of the graph, i.e. the number of nodes, ${\lvert{\mathcal{V}}\rvert}$. Thus, for large networks it can be assumed that the bound is significantly lower than the actual number of nodes: